Minutes of the 13th Committee Meeting of the Dublin University Rifle Club

Moderator Steven Watterson Minutes taker Mark Dennehy Present Steven Watterson (Captain),

Benjamin Jones (Treasurer),

Mark Dennehy (Secretary/Match Secretary), John Keeney (Ammo&Equipment Officer), Alan Smyth (Chief Range Officer), Conor McDermottroe (Public Relations Officer), Deirdre Ninaber (Ladies’ Coordinator), Peter Byrne

Absent(with apologies) Absent Jason Sibley

Date 6 Nov 2001 1850 – 2210 Distribution DURC Committee

Overview of topics

1 Peter Byrne 1

2 Petrol 1

3 Formality 2

4 Targets 2

5 Secretary’s Report 2

6 CRO Report 2

7 A&E Report 2

8 Capital Grant 2

1 Peter Byrne

8.1 Aluminium stocks for old .22 rifles . . . . . . . . . . . . . . . . . . . . 3 8.2 Shooting Trousers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 8.3 Electric target changers . . . . . . . . . . . . . . . . . . . . . . . . . . 4 8.4 Shooting Boots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 8.5 Air Rifles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 8.6 Gun Safe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 8.7 Scopes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 8.8 Jackets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 8.9 Aluminium Buttplate/Hook assemblies . . . . . . . . . . . . . . . . . . 8 8.10 Shortlist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

9 DURC 10m Air Rifle Open 9 9.1 Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 9.2 A&E . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 9.3 IT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 9.4 Rotas . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 9.5 Costs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 9.6 Advertising . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

10 AOB 10

1 Peter Byrne

It was decided by the entire committee (except for Jason who was absent) to co-opt Peter Byrne onto the committee due to his continuing committment to the club and his presence at previous informal committee meetings. Peter accepted as graciously as was possible while kicking and screaming.

2 Petrol

Ben wanted people to turn in receipts for petrol from the UCD and RRPC weekends.

3 Formality

Steven pointed out that the last week’s formality had gotten a lot done in a shorter space of time and proposed we continue to use it. Noone objected.

4 Targets

4 Targets

Steven pointed out that the current NSRA targets the club has have a problem in that they tear badly due to their thinness. Mark happily said “I told you so”. Buying Edelmann targets was accepted as the course of action. John reported not having prices yet and that he would get some the following day. Single-bull targets were mentioned and Mark was adamant that we could not buy them from Dave Cooney as the ones used in the Nationals had come from his storage area and were so damp that the Nationals’ Finals had been a shambles with many challanges. It was agreed by all that it was a priority to get targets soon.

5 Secretary’s Report

Mark reported that AHG and FWB had replied to him regarding purchasing capital gear and that AHG had sent catalogs and were offering a 10% discount on AHG items and 7.5% on JHG items. FWB, however, were asking that we talk to Dave Cooney as the Irish distributor. It was not felt that this would give a good discount.

6 CRO Report

Alan reported having talked to Liam about his missing RO duty slots. Liam apologised but said he thought Steven had him covered for those nights. Steven said he should have arranged something on those nights.

7 A&E Report

John gave his impressions of the new DUCAC treasurer which were not favorable. She did not, in his opinion, seem to understand that our jackets were not personal gear and she had expressed her policy of not paying for personal gear. John was adamant that we must stress that jackets, etc, are not personal gear for this reason.

Mark at this point mentioned that Eric had suggested that ESB would be willing to sponsor us to buy some jackets, but that nothing had gone past the possible stage yet.

8 Capital Grant

John and Ben were asked for their impressions of the DUCAC grant decision-making process from past experience as treasurers. John replied that with a new Treasurer in DUCAC, past experience was not terribly informative. Ben reported that last years grant came to £1300, approximately 1650. He expected us to receive a similar amount this year. Ben also pointed out that we fufilled most if not all of DUCAC’s criteria for getting large grants.

8 Capital Grant

A list of options was taken by Steven:

Aluminium stocks for old .22 rifles  Shooting Trousers
Electric target changers  Shooting Boots
Air Rifles (PCP?)  Gun Safe
Junior Air Rifles  Jackets

Hook/Buttplate assemblys for old .22 rifles

Discussion then followed on each item.

8.1 Aluminium stocks for old .22 rifles

Mark proposed this option because:

1. The aluminium stocks were aestheticly pleasing and that the look of them alone raised interest in shooting to a higher level if that would allow people to use them

2. It would allow us to reuse old target rifles with perfectly acceptable actions/barrels but which were not used today due to old stocks, e.g. #3

3. It would allow us to field more 3P shooters for competition

Discussion was then opened to the floor.

Alan pointed out that we had two 3P rifles allready and they weren’t used for 3P. He also felt that #3, being a standard rifle, was past it’s point of usefulness anyway. Conor wondered about the cost of this option. Steven suggested about £800-900 from the NSRA. Alan had a problem with the elitism that the stocks would promote. Mark reported that AHG’s price was 989 for a 2213 stock, and the NSRA price was £754. Alan commented that he felt the level of adjustment on the stocks was a good idea, especially as he had recently been told to cut a few inches off his own stock. John pointed out that the stocks were good for prone shooting as well as 3P. Steven felt that the cost was quite high and that setting an alu stock up was quite tricky and involved. Alan suggested waiting for the new range as the equipment reuse argument would be far more weighty then. Peter suggested that alternatives would be a better use of money.

Steven called for a vote to keep alu stocks on the shortlist.

For  Against  Abstention
2  3  3

Motion rejected

8.2 Shooting Trousers

Steven proposed this option because:

1. They would add to all the Air Rifle scores by up to 40 or 50 points.

2. He felt they contributed more to good shooting in Air than the jackets did.

3. The sharing trousers problem would be negated by wearing jeans underneath them.

Discussion was then opened to the floor.

John wondered if we could wear jeans under the trousers. Steven said that this was not a problem - Mark said that leggings were normally worn and that jeans wouldn’t be a problem. John’s comments on lycra shorts were not recorded for posterity. John wondered if we could get enough sizes to adaquately cover the range of people in the club. This worried Conor also. Alan wondered about the administration of selecting trousers for people, preventing gear conflicts and so forth. Mark found a price in the AHG catalog of 121. Ben wondered if the length of the trousers mattered. Steven responded with a quick explaination of how the trousers operated, mainly pointing out that a variation in waist size up to 1 or 2 inches was fine and that the only part of the leg that counted was from thigh to waist. Deirdre wondered if adjustable waists were available, Steven replied that Marks were an example of this. Peter thought they were a good idea to get air scores up. Mark reminded everyone to budget for suspenders to prevent serious levels of wear and tear on the cuffs and zips of the trousers. Ben wondered if there was time in our half-hour details to get into and out of the trousers. Ben also thought that their price was low enough that they might be purchasable on the Non-Capital grant. John pointed out a rumour from the UIT email list that the ISSF was planning on changing regulations for the thickness of jackets and trousers and that this might result in them being worthless in a few years. Steven gave a brief account of the ISSF rulechange being discussed in an ISSF technical committee and it was pointed out that such rules would take time to reach Irish national competition levels. Mark also pointed out the many unconfirmed rumours in this area to show we shouldn’t take them seriously until they are actually written down.

Steven called for a vote to keep trousers on the shortlist.

For  Against  Abstention
8  0  0

Motion passed

8.3 Electric target changers

Steven wanted to know how broken our current electric changers were, but noone knew for certain, though all felt that repair should be within our capabilities. Prices were unknown but were thought to be relatively cheap. John agreed to check our current changers.

Steven called for a vote to keep target changers on the shortlist.

For  Against  Abstention
0  8  0

Motion rejected

8.4 Shooting Boots

Mark proposed this option because:

1. Shooting in bare feet was the best option for scores but put enormous stress on the achilles’ tendon, causing pain and possible chronic damage.

2. Shooting in boots gave a better score.

Discussion was then opened to the floor.

Alan said he’d never noticed boots adding to his scores before, but he did feel more comfortable in boots. John thought people should be expected to have a pair of normal shoes with low heels and flattish soles as part of their personal air rifle kit, and that we shouldn’t be buying them for our shooters. Peter said he’d never noticed any improvement in his scores from using boots. Mark said that it didn’t help achieve a score so much as allow you to achieve it for longer by increasing your comfort level. Ben wondered how many pairs we would need. Peter felt we’d need one of every size from 12 to 3. Alan felt it was too personal a piece of gear to share with 400 other people. Mark found a price of 99 from AHG for basic boots. John felt that this would make the problem of access to a specific set of club gear worse. Summary points were called for by Steven. Mark felt they had medical and score benefits. Alan felt they were too expensive. Peter felt trousers gave more points for roughly the same amount of money. John felt they were personal kit. Steven and Mark both felt this was not true.

Steven called for a vote to keep boots on the shortlist.

For  Against  Abstention
2  3  3

Motion rejected

Steven was asked for his personal opinion - he reported he had never shot without boots.

8.5 Air Rifles

Ben proposed this option because:

1. We don’t have enough air rifles. We had to borrow two at the UCD open. We have many .22 rifles but only 6 air rifles, of which one is being repaired in germany.

Discussion was then opened to the floor.

Alan felt our current 601s had seen too much wear and tear and new rifles were needed to step in. John said we needed more variety in our rifles to be able to match our air shooters to rifles better. Deirdre thought we should get a junior rifle for girls because the balance was too akward for many on the full-size rifles. Ben wondered if DUCAC would pay for Junior’s equipment. Conor pointed out that it was a size, not an age related distinction, and that up to an eighth of our people in Freshers’ week had had problems due to the size of the rifles, or difficulty in cocking them. Steven pointed out that girls should be shooting with a chair to rest their rifles on for leverage. Mark suggested lefthanded rifles would be a good investment. Ben replied that only 10% of the population was left-handed (Mark countered that 40% was a more likely figure). John pointed out that the .22 left-handed rifle was so delapidated that if we were to buy a left-handed rifle, we should start with a .22. Alan pointed out that right-handed shooters still had gear conflicts and they had to cope with it. It was wondered how serious a problem it was.

John proposed that we get a PCP air rifle. He noted that when he used Ray’s P70 his score went up by 30 points because the muzzle velocity was higher, thus reducing the importance of the follow-through. Alan pointed out that PCP rifles would eliminate the aerobics workout that charging was during shooting. Mark gave testimonial on his new PCP rifle and fully backed the idea as it raised scores, allowed easier loading for less muscular shooters and allowed easier 3P shooting. Steven wondered where we’d get the air compressed. Mark replied that he had talked to the scuba club about his own rifle and they were amenable to letting us fill up there, but it was pointed out by Ben and John that getting our own cylinder was a better idea. Safety was discussed - a five-year hydrotesting would be required. Steven remained unconvinced the extra hassle was worth the effort, and didn’t think that extra points would show up in competition due to other factors causing stress. Peter pointed out that his last string in the UCD open had been poor as he was exhausted from charging the rifle. John pointed out Ray’s back problems caused by not breaking position during a match. Mark replied that Ray’s age was a factor in those problems. Conor wondered about if we had enough safe space.

Steven called for a vote to keep air rifles on the shortlist.

For  Against  Abstention
8  0  0

Motion passed

8.6 Gun Safe

Steven pointed out that the safe was currently over-full and we had rifles in Front Arch. Conor pointed out that we had rifles in the club safe held by individuals - Mark, Steven, Ray and Keith. Conor also felt having any rifles in front arch with bolts and ammo was a serious problem, especially with security moving out of there soon. Ben wondered where we’d put the safe. Steven said we would find the space, that there was plenty on the range. John said we had to have it in a lockable area. John also said not to bring it up with security as a security problem. Mark suggested that people with personal rifles chip together to buy a safe for those rifles on the range. John suggested we just buy a big safe and those with personal rifles chip in. Peter wondered if we should trade our current safe in - this was felt to be a bad idea as ours is a high-quality safe. John and Mark discussed the possibility of second-hand safes which wandered to a discussion on the outcome of the NARGC gun safes case. A discussion on prices took place and it was felt that £150 upwards from Geoff Cooney was about the best that could be done.

Steven called for a vote to keep a gun safe on the shortlist.

For  Against  Abstention
4  1  3

Motion passed

8.7 Scopes

John and Alan commented on the current scopes. They were in pretty poor condition, with only 4-6 being usable, none of which were usable at 100 yards. A better breakdown was listed as:

25yards 7/8 servicable scopes
50m  4/5 servicable scopes
100yards 0 servicable scopes

Though these figures were contested by Mark, who said that the current scopes caused eyestrain and only one or two were adaquate. Steven wondered if they could be repaired and if not, what would the replacement cost be. Alan said that repairs had proved temporary in the past. John felt they couldn’t be repaired but should be cleaned professionally. Mark felt that they couldn’t be repaired as the optics were simply not up to the task. John pointed out that we couldn’t afford a decent scope. Mark found prices - 250 and 350. Steven wondered if the stands were damaged - Mark and Alan agreed that they were repairable. John pointed out that he had examined the 250 model scope and not found it to be any better than our current ones in practise. Alan pointed out that the scope John was comparing it to was one of our better scopes. Mark pointed out that the 350 model was the same as Rhona Barrys, and that that scope was excellent. John pointed out that that model did not fit our stands, and neither did the other model. Ben felt we had enough scopes allready. Mark felt the eyestrain problem was being understated. Steven felt it was not the biggest problem. John said he’d never had eyestrain problems before. Alan felt that overall our scopes were of poor quality. Conor said that they gave a poor sight picture. Mark suggested replacement eyepieces might by a useful idea. Ben wondered if they would help. Mark found a price of 60 for them. John and Mark commented on their differing opinions on JK Walsh’s scope. Steven called for pros and cons. Peter felt that we needed a scope but not for 100yards which we only shot once or thrice a year. Mark felt we needed one due to the eyestrain and sight pictures from our current fleet. However he wondered if eyepieces were an acceptable comprimise. Alan pointed out that the sighting-in phase is highly important and a good scope a necessity. Deirdre had nothing to add. Conor agreed with Mark. Ben felt strongly that we allready had enough scopes and more would be a waste, especially as we only shot 100yds three times a year. John said it was a lot of money for little return and maintainance was a better solution. Peter pointed out that even though we don’t shoot 100yds often, we still had to be able to shoot it. Mark asked Ben if eyepieces could go on the NonCap grant, Ben replied yes.

Steven called for a vote to keep scopes on the shortlist.

For  Against  Abstention
4  2  2

Motion passed

8.8 Jackets

Mark pointed out that there had been a few people that didn’t easily fit any jacket this year, however this topic was dismissed without much discussion as it was felt not to be needed at present.

8.9 Aluminium Buttplate/Hook assemblies

John advocated this choice. Mark found a price of 200 for the AHG and MEC buttplates, but reminded everyone that this was not including a buttplate carrier. Steven wondered if they were useful, Mark, John, Peter and Alan said it was. Mark wondered if Air Rifle buttplates could be ordered. This was dismissed. Steven wondered if we would upgrade all our rifles - John replied that no, only one or two would be done. John pointed out though, that retrofitting was not easy or cheap, or even necessarily possible.

Steven called for a vote to keep buttplates on the shortlist.

For  Against  Abstention
8  0  0

Motion passed

8.10 Shortlist

Steven then listed the shortlist:

• Shooting Trousers ( 150)

• Air Rifle ( 1200-1300)

• Gun Safe ( 300)

• Scope ( 300)

• Buttplates ( 300)

Ben noted that we could apply for all and fall within a reasonable budget. A discussion on numbers of items ensued. Peter noted that he had intended to train that night and the committee’s response was “welcome to the committee, peter!”.

Alan wondered if we could allocate priorities in the grant application. A discussion on how many trousers were needed cropped up, with both ben’s estimate of 2 and steven’s of 5 being seen as unrealistic. Money-back guarantees were discussed, and if the gear

9 DURC 10m Air Rifle Open

is not custom-made, it should be possible to return it. On the question of who should be using the trousers (and how many were therefore needed), Steven remarked that “it doesn’t take skill to use trousers properly”. He offered to demonstrate by using either his trousers or marks, but the offer was declined as soon as the committee stopped laughing. After some negotiation, the following was settled on for the application:

3 Shooting Trousers
1 Air Rifle
1 Gun Safe
1 Scope
1 Buttplate assembly
Total Cost 2550

This was accepted by all, but Alan registered his unhappiness at the amount of money being spent on the Air Rifle.

9 DURC 10m Air Rifle Open

Mark gave a list of sponsors asked to help and reported that Meteor had turned us down. Alan asked not to be sent the emails regarding the sponsorship and people in general said they’d prefer if they only got the emails from sponsors agreeing to help.

9.1 Format

2 hours per detail, beginning Saturday at 1100, Sunday at 1000, with a 10-shot final at 1630 on Sunday.

9.2 A&E

Graduated guages and targets are needed. John will acquire these as soon as possible.

9.3 IT

Mark reviewed last year’s setup. Steven was worried it would be a lot of work, but most people involved last year disagreed. John suggested that we take a modular approach instead of an all-or-nothing one.

9.4 Rotas

Ben wondered if a shooter could score his own cards. Steven offered to disregard his own scores, but was overruled by Mark and John as he was likely to put in the best score from DURC. Mark said a pool of scorers was allowed in ISSF rules. John said Keith shouldn’t do it all this year again. Steven and John pointed out that we should be expecting no time off that weekend (with the exception of Conor and Peter who worked weekends. Mark however, promised Conor cake and song if he came down on the 24th which was Conors birthday). Mark, John and Steven indicated they were happy to do

10 AOB

RO duty and stats duty on the weekend. Alan said he’d draw up the rotas. It was agreed to have 2 ROs, 1 scorer and 1 data entry operator at all times.

9.5 Costs

The event, it was agreed, must break even or make a profit. Steven wondered about UCD range hire - Alan felt £30 would be a likely maximum, and free might be possible. Targets were estimated at £200 to £300 by John. Ben did some estimates and came up with us raising £300 in fees. A debate on the fees level arose and went on for a while - it was felt that UCD and WTSC having charged £9 we couldn’t charge £10 without having good prizes. Alan felt DUCAC shouldn’t pay for targets, Mark strongly disagreed. Ben reminded people of the £1 NTSA levy. After some discussion with Ben using his little whiteboard to full effect and Steven at one point suggesting an orgy in the UCD range to increase attendance, it was agreed to charge £9 for adults, £6 for students/DURC members/Juniors. This left a budget of £150 for prizes.

9.6 Advertising

Mark said he was going to take two days to contact potential sponsors directly and then he would print the flyers for the open and post them to all concerned. More would be printed and handed out at the WAIDAM open that weekend.

10 AOB

Peter wanted shorter meetings. Everyone agreed, but noone had a viable solution.

Mark pointed out that many people had asked about getting their own gear and that we should invite them to order their items when we made our order to avail of our discount. This was agreed to be a good idea and that we would do so.

T.B.N.O.B.T.M.W.D.C.

Steven Watterson, Club Captain

Mark Dennehy, Club Secretary

10 AOB

Tasks agreed on at meeting